Q3-Evaluate the integral: dx Q4-Evaluate the integral: tan x dx (x + 1)
Answer #1

(3): Let I I= n+3 dn 2+8x+7 こ (n+3) au **+74+x+7 IS (x+32 du x+3) dn *(X+7)+1(n+7) (n+1)(x+7) bartial fraction: using x+3 + BEtk = 1 / 2 1 4 4 + 2 het L+8+zu nt3 du 2 tup 2 x+84 +7 3 3 +7 SA Intel + 2/5 Sam Ian |x+71+c N C Inlla+11+ 2ln(a+7)+c (4) Le